OpenStudy (anonymous):

An oxide of phosphorus contains 56.4% phosphorus and 43.6% oxygen. It's relative molecular mass is 220. Find both the empirical and the molecular formula of the oxide'

OpenStudy (anonymous):

moles of P = 56,4g/30,974g/mole = 1,82 moles P moles of O = 43,6 /15,999 = 2,73 moles of O Converting to the simplest ratio: For P: 1,82/1, 82 =1 For O: 2,73/1,82 = 1,5 which I guess approximates 1 P and 2 Oxygens. PO2 -> the empirical formula I suppose this is how it's done, but Im not quite sure.